OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below. Show your work. (3 points) x^2 – 36 = 0 and x^2 = 8x – 12 Part 2: Describe what the solution(s) represent to the graph of each. (2 points) Part 3: How are the graphs alike? How are they different? (2 points)

OpenStudy (anonymous):

Is the formula to these -b +- (sqrt) b^2 - 4ac / 2a?

OpenStudy (campbell_st):

\[x^2 - 36 = 0=====> (x -6)(x+6) = 0\] x = 6 or -6 number 2 rewrite the equation x^2 - 8x + 12 = 0 this also factorises (x - 6)(x - 2) = 0 the solutions are x =6 and x = 2

OpenStudy (lgbasallote):

@careless850 you can use that formula too

OpenStudy (campbell_st):

the solutions to the equations are where the graphs cut the x -axis y = x^2 - 36 cuts the x-axis at -6, 6 is symmetric about the y axis has a y intercept of y = 36 and is concave up. y x^2 - 8x + 12 cuts the x=axis at x = 2 and 6 , is concave up, has a line of symmetry of x = 4 cuts the y axis at y =12 and has a minimum value of y = -4

OpenStudy (anonymous):

So confusing dude...

OpenStudy (campbell_st):

in which part

OpenStudy (anonymous):

Is there a formula I can use?

OpenStudy (anonymous):

A simple algebraic formula..

OpenStudy (lgbasallote):

since you know the quadratic formula you can apply it in x^2 - 36 a = 1; b= 0; c = -36 \[\LARGE x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] \[\LARGE x = \frac{ 0^2 \pm \sqrt{0^2 - 4(1)(-36)}}{2(1)}\] got that?

OpenStudy (campbell_st):

you can use \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] just identify a b and c in each equation

OpenStudy (campbell_st):

there is no simple algebra formula for solving quadratics... factorising or using the formula shown by labasallote

OpenStudy (anonymous):

That one

OpenStudy (anonymous):

You just did.

OpenStudy (campbell_st):

well thats the only one... but you need to remember you need evaluate the formula fully if the examiners are looking for solutions

OpenStudy (anonymous):

Okay so can we try the first one with that formula? step by step?

OpenStudy (campbell_st):

ok.... what is a= b= and c =

OpenStudy (anonymous):

a is 1 b is 1 and c is -36.

OpenStudy (campbell_st):

there is no b in x^2 - 36... only a and c

OpenStudy (campbell_st):

or more correctly b = 0

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

So, a is 1 b is 0 and c is -36.

OpenStudy (campbell_st):

yes... good start

OpenStudy (anonymous):

Now what do I do from here?

OpenStudy (campbell_st):

substitute them into the formula write it on the drawing pad

OpenStudy (anonymous):